OpenStudy (anonymous):

Help me make a equation for this chart? Please! x=1 y=1 x=2 y=1+2 x=3 y=1+2+4 x=4 y=1+2+4+8 x=5 y=1+2+4+8+16 x=6 y=1+2+4+8+16+32 x=7 y=1+2+4+8+16+32+62

OpenStudy (anonymous):

That last one is supposed to be " .. 16+32+64" right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you still need help? It says the question is closed.

OpenStudy (anonymous):

Yes, I closed it by mistake.

OpenStudy (anonymous):

Ok, do you know the general formula for numbers that look like those? 1, 2, 4, 8, 16, etc.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ok. They're powers of two. \(2^0 = 1\) \(2^1 = 2\) \(2^2 = 4\) \(2^3 = 8\) \(2^4 = 16\) and so on...

OpenStudy (anonymous):

oh yes I know those sorry I wasn't quite sure what you were talking about.

OpenStudy (anonymous):

Ok! Well, if you know that, then think about the series you have. Pick a couple of them at random and add them up and see if the result looks familiar.

OpenStudy (anonymous):

Better yet add one of the larger ones up by hand so you can see the sum at each phase and really think about what you're getting...!